OpenStudy (anonymous):

Karen bakes cookies. She has observed that the number of cookies that turn out well in each batch varies with the size of the batch and is given by the function f(x) = 2x – 23, where x is the size of the batch (number of cookies in each batch). She never bakes fewer than 12 cookies in a batch. If the size of five batches is {20, 25, 26, 27, 34}, how many cookies turn out well in each batch?

OpenStudy (anonymous):

f(20) = 2(20) - 23 = 17 f(25) = 2(25) - 23 = 27 f(26) = 2(26) - 23 = 29 f(27) = 2(27) - 23 = 31 f(34) = 2(34) - 23 = 45 you plug in whatever your x value is when you see f(x) and then evaluate your equation.

OpenStudy (anonymous):

so f(x) basically means y, where y is the value you want. if this were to be written in y = form, it would be: y = 2(20) - 23 y = 17

OpenStudy (anonymous):

does that make sense? @zxcvbnm08765